The interest rate for an annuity  Personal Finance Problem   Anna Waldheim was seriously injured in an industrial accident. She sued the responsible parties and was awarded a judgment of

$1,500,000. Today, she and her attorney are attending a settlement conference with the defendants. The defendants have made an initial offer of $119,519 per year for 24 years. Anna plans to counteroffer at $192,695 per year for 24 years. Both the offer and the counteroffer have a present value of $1,500,000​, the amount of the judgment. Both assume payments at the end of each year.

a.  What interest rate assumption have the defendants used in their offer​ (rounded to the nearest whole​ percent)?

b.  What interest rate assumption have Anna and her lawyer used in their counteroffer​ (rounded to the nearest whole​ percent)?

c.  Anna is willing to settle for an annuity that carries an interest rate assumption of 99​%. What annual payment would be acceptable to​ her?
Answer #1

1.
=RATE(24,-119519,1500000)
=6.0000%

2.
=RATE(24,-192695,1500000)
=12.0000%

3.
=PMT(9%,24,-1500000)
=154533.84